This morning, you borrowed $150,000 to buy a house. The mortgage rate is 7 percent. The amortized loan requires you to pay $10,000 in principal plus interest each year over 15 years. The first payment is due one year from today. How much is the second payment?
Answer #1

Loan amount = $150000

mortgage rate = 7%

1 payment = $10000 + interest

interest on loan amount = loan amount* interest = 150000*0.07 = $10500

So, 1st payment = 10000+10500 = $20500

principal amount remaining = 150000-10000 = $140000

for 2nd payment interest = 140000*0.07 = $9800

So total 2nd payment = 10000+9800 = $19800